OpenStudy (anonymous):

-{2a+[3a-(2b-6)+3(b-4)]-2[a-(3b+1)]} simplify..

OpenStudy (anonymous):

help

OpenStudy (anonymous):

probably =-{2a + [3a -2b+6+3b-12]-2a+2(3b+1)} = -{3a+b-6+6b+2} =-{3a+7b-4}= -3a-7b+4

OpenStudy (anonymous):

angela can u show me ur solutions?

OpenStudy (anonymous):

because my ans. is -3a-7b+4

OpenStudy (anonymous):

its (-3a-7b+4)

OpenStudy (angela210793):

ur right i have added -6 instead of 6...sorry :)